OpenStudy (anonymous):

For the following system, if you isolated x in the first equation to use the Substitution Method, what expression would you substitute into the second equation? -x - 2y = -4 3x + y = 12

OpenStudy (anonymous):

these are the possible answers... a. -2y - 4 s. 2y - 4 c. 2y + 4 d. -2y + 4

OpenStudy (anonymous):

x=4-2y in the first equation substitute into the second equation 3*(4-2y)+y=12 12-6y+y=12 -5y=0 y=0 x=4

OpenStudy (amistre64):

solving an equation for a given variable is like chapter 1; this is like chapter 12 ....